obtain a relation between the period of a simple harmonic motion and acceleration at unit displacement
1
Expert's answer
2021-09-12T18:58:59-0400

The displacement for a simple harmonic motion

"x(t)=x_0\cos\omega t=x_0\cos\frac{2\pi t}{T}"

The acceleration

"a(t)=x''(t)=-\frac{4\pi^2}{T^2}x_0\cos\frac{2\pi t}{T}=-\frac{4\pi^2}{T^2}x(t)"

So

"T=2\pi\sqrt{x/a}"